Let
W = ((X1, X2, X3) belongs to R³ : X2 + X3 = 0).
Show that W is a subspace of R³ . Find two
subspaces W1 and W2 of R3 such that
R³ = W (direct sum )W1 and R³ = W (direct sum) W2 but W1 (not equal to) W2.
1
Expert's answer
2020-11-18T19:27:07-0500

Given

"W=\{(x_1,x_2,x_3)\R^3:x_2+x_3=0\}"

Let "a,b\in \R\&(x_1,x_2,x_3),(y_1,y_2,y_3)\in\R^3"

Now,


"a(x_1,x_2,x_3)+b(y_1,y_2,y_3)=(ax_1+by_1,ax_2+by_2,ax_3+by_3)"

Note that

"ax_2+by_2+ax_3+by_3=a(x_2+x_3)+b(y_2+y_3)=0\\
\implies (ax_1+by_1,ax_2+by_2,ax_3+by_3)\in W"

Hence, W is a subspace of "\R^3" .


Let us define a subspace

"W_1=\{(x_2,x_2,0)\}\\
W_2=\{(x_2,2x_2,0)\}"

Clearly,


"(\star)\,W_1\cap W_2=\{0\},W_1\cap W=\{0\},W_2\cap W=\{0\}"

And,


"\R^3=W\oplus W_1\\"

As, for any "\vec v\in \R^3" we can write "\vec v=\vec u+\vec w" for some "\vec u\in W,\vec w\in W_1" and hold "(\star)" .


Similarly for

"\R^3=W\oplus W_2"